Two cars pass on a highway while traveling in oppisite directions.One car is traveling 6 mph faster than the other car.After 1 hour and 30 minutes the two cars are 171 miles apart.Find speed of car.
Answer by mananth(16949)

Car I x mph
Car II x + 6 mph
they are moving away from each other
so add up the speeds
x + x + 6
= 2 x + 6
t= d/r
t= 1.5 hours
d=171
r= 2 x + 6
1.5 =171/( 2 x +6)
1.5 ( 2 x+ 6 )= 171
3 x + 9 = 171
3 x = 171 - 9
3 x = 162
/ 3
x= 162 / 3
x= 54 mph

Car I = 54 mph
Car II= 54 + 6 = 60 mph
m.ananth@hotmail.ca